i know the answer to the word problem i am presenting, what i want to know is, is there a mathematical way of solving it (i used trail and error)?
a car travelling at 90 km/h is 500 m behind another car travelling at 70km/h in the same direction. how long will it take the 1st car to overtake the second? the answer is 1.5 minutes

Answer by TimothyLamb(4379)

---
relative speed:
90 - 70 = 20 km/hr
---
s = d / t
t = d / s
---
t = 500 m / (20 km/hr * 1000 m/km * 1/60 hr/min)
t = 1.5 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php